in right angle triangle the sides of right angles are 3 and 4 find its hypoteneous?
Answer by stanbon(75887)

in right angle triangle the sides of right angles are 3 and 4 find its hypoteneous?
------
Equation:
h^2 = 3^2 + 4^2
---
h = sqrt[9 + 16]
---
h = sqrt(25)
---
h = 5
===============
Cheers,
Stan H.
===========